OpenStudy (anonymous):

for 0x=90 cosx=1/3 --> x=70.53 what're the other roots?

hartnn (hartnn):

cosx=1 -->x=90 ???

hartnn (hartnn):

you took y =cos x y=1 cos x = 1 for which angle ?

OpenStudy (anonymous):

x=0,360

hartnn (hartnn):

those are not in the range, right ? so no roots for cos x = 1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cosx=1/3 x=70.528

hartnn (hartnn):

\(\cos (360-x)=\cos x\)

OpenStudy (anonymous):

289.4

hartnn (hartnn):

yes, correct :)

OpenStudy (anonymous):

just 2?

hartnn (hartnn):

yes.